OpenStudy (anonymous):

The equation of line AB is (y−3) = 5 (x − 4). What is the slope of a line perpendicular to line AB?

OpenStudy (anonymous):

The slope is the reciprocal of 5 which is -1/5

OpenStudy (the_fizicx99):

Solve the equation first, (y−3) = 5 (x − 4) Apply distributive property to 5(x - 4) (y−3) = 5 (x − 4) (y−3) = 5x - 20 Now add both sides 3 to isolate y on the left side, y-3 - 5x - 20 +3 +3 y = 5x -17 When the slope is perpendicular its the reciprocal , so you flip it and multiply it by -1 5 = 1/5 * -1 = -1/5 So y = -1/5x - 17

OpenStudy (anonymous):

Thankyou for explaining

OpenStudy (the_fizicx99):

yw